OpenStudy (anonymous):

find two pairs of polar coordinates for points(1,1) satisfying -2pi

OpenStudy (anonymous):



OpenStudy (anonymous):

i want to find the polar coordinates

OpenStudy (anonymous):

the polar coordinates are \((r,\theta)\)

OpenStudy (anonymous):

yes i know how to find r and theta but then the condition -2pi<theta<=0 is to be fullfilled to find polar coordinates??

OpenStudy (anonymous):

\[r=\sqrt{x^2+y^2}\] \[\theta = \tan^{-1} \left( \frac {y}{x} \right)\] but theta must be adjusted to fit your criteria, i.e., \(-2\pi<\theta \le 0\)

OpenStudy (anonymous):

look at the point. what would theta be regularly? that is, if 0 < theta < pi/2

OpenStudy (anonymous):

as (1,1) is in ist coordinaye should we add and subtract 90 to thete ??

OpenStudy (anonymous):